OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 16 times y cubed − the square root of y cubed 3 times y times the square root of y 3 times y squared times the square root of y 4 times y squared times the square root of y 4 times y times the square root of y

OpenStudy (anonymous):

@JimmyC @phi ???

OpenStudy (anonymous):

How many times do I have to tell you to not ask for math. I'm taking summer school for math right now :P

OpenStudy (anonymous):

@JimmyC sorry lol

OpenStudy (phi):

First, write out the problem in "math" \[ \sqrt{16 y^3} - \sqrt{y^3} \] look for "pairs" inside the square root, and then use the rule \[ \sqrt{ x\cdot x} = x \sqrt{1} = x \] also, remember the little number in the upper right means multiply by itself that many times. The idea is : write numbers as factor of prime numbers and write out the letters to a power in expanded form. then look for pairs: \[ \sqrt{16 y^3} - \sqrt{y^3} \\ \sqrt{2\cdot 2 \cdot 2 \cdot 2 \cdot y\cdot y\cdot y}- \sqrt{y\cdot y\cdot y} \] can you do the next step: pull out pairs?

OpenStudy (anonymous):

you pull out the y^3? all of them?

OpenStudy (phi):

only pairs. anything left over stays inside the square root. and when you pull out a pair you put only one of the pair on the outside